OpenStudy (ballerlea):

What are the x- and y- intercepts of the following equation: y = -0.05x + 16

OpenStudy (luigi0210):

X-intercept is when y=0 Y-intercept is when x=0.

OpenStudy (tkhunny):

And, just for the record. y = -0.05x + 16 y = -x/20 + 16 x/20 + y = 16 x/320 + y/16 = 1 There be secrets in there!